# suppSupplementary References

### v2

@@ 4.2.1 A weak stochastic order based on the Cramér distance
[…]
for all $10.5<\tau,\xi<$ 1. To the best of our knowledge, the weak stochastic order is a new order relation, and we discuss its properties after establishing its equivalence with the shift components of the $\operatorname{CD}$.
@@ 4.2.2 A relaxed stochastic order based on Wasserstein distances
for all $10.5<\tau<$ 1. To the best of our knowledge, the relaxed stochastic order relation is new to the literature, and we briefly discuss its properties at the end of this section.
@@
Let $F$ and $G$ be probability distributions with continuous quantile functions $F^{-1}$ and $G^{-1}$, respectively. Then, for all $p\in\mathbb{N}$,
@@ 6 Discussion
As a consequence of our interval-based approach, the proposed decompositions attribute the distance between two distributions $F$ and $G$ entirely to differences in location and dispersion. Of course, distributions are often characterized by additional (higher-order) properties such as skewness or kurtosis attributed to their shape. In our decompositions, differences in higher-order properties sometimes lead to both shift or dispersion components being nonzero simultaneously (e.g., in Example 2.2), but there are no clear-cut connections. The nonparametric nature of our decompositions that aggregate fundamental comparisons of central intervals does not attest to differences in shape, while simple decompositions based on comparing summary statistics, which may do so, provide a rather superficial comparison and are not available for most of the studied distances. Accommodating additional shape components likely requires considerably revised techniques and is reserved for future work.
Finally, the decompositions are thus far descriptive tools that appear to be promising both from a theoretical and applied perspective. A next step will be to develop inference techniques that assess the statistical significance of the components in standard one- and two-sample settings as well as more involved settings such as aggregated inflation expectations. While resampling techniques such as permutation tests or the bootstrap are well suited to assess the statistical significance of the overall distance, assessing the significance of individual components is challenging. As the null hypotheses of no difference in shift or location are complex composite hypotheses, simulating them does not seem feasible. A careful look at the sum of the dispersion components of the area validation metric reveals that it can be rewritten as a sum of conditional expectations, which may well facilitate inference for this particular component. We view such developments as a natural avenue for future research.
@@ Acknowledgements
We thank Thomas Eife and Thordis Thorarinsdottir for sharing the data sets for the two applications and Christoph Becker, Andreas Eberl, Thomas Eife, Tilmann Gneiting and Melanie Schienle for their insightful comments. T. Dimitriadis and J. Resin (project 502572912) as well as J. Bracher (project 512483310) gratefully acknowledge support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). D. Wolffram was supported by the Helmholtz Association under the joint research school HIDSS4Health – Helmholtz Information and Data Science School for Health. All authors acknowledge the Heidelberg Institute for Theoretical Studies (HITS) as secondary affiliation.
@@
The supplement contains proofs and derivations together with additional illustrations, counterexamples, and details on approximations. Section S1 graphically illustrates the CD decomposition for a range of $\beta$ values. We provide closed-form expressions for the $\operatorname{WD}_{p}$ decomposition in Section S2. Section S3 provides additional counterexamples. Section S4 provides a general purpose approximation of the CD decomposition and gives details on the approximations used in Section 5.2. Finally, Sections S5S7 contain all proofs of the results presented in Sections 24, respectively.
@@
Consider the distributions $F$ and $G$ that are illustrated in panel (a) of Figure S2. The distribution $F$ places $\frac{1}{4}$ probability mass at $-1$ and $\frac{3}{4}$ probability mass at 0 and $G$ is given by the location-scale transformation $G(x)=F(\frac{x-0.5}{1.8})$. The first three $p$-th power $p$-Wasserstein distances between $F$ and $G$ and their decompositions are
@@
S4 Approximation
In many cases, such as in the application in Section 5.2, the CDFs $F$ and $G$ under consideration are only partially known at a finite number of values or quantiles (that can even differ between $F$ and $G$). Here, we outline two approximation strategies for such scenarios.
S4.1 Distributions in a quantile format

### v3

@@ 4.2.1 A weak stochastic order based on the Cramér distance
$\max\big{\{}F^{-1}(\tau)-G^{-1}(\xi),\,F^{-1}(1-\tau)-G^{-1}(1-\xi)\big{\}}\;\geq\;0$ (27)
for all $0.5<\tau,\xi<1$. To the best of our knowledge, the weak stochastic order is a new order relation, and we discuss its properties after establishing its equivalence with the shift components of the $\operatorname{CD}$.
@@ 4.2.2 A relaxed stochastic order based on Wasserstein distances
for all $0.5<\tau<1$. To the best of our knowledge, the relaxed stochastic order relation is new to the literature, and we briefly discuss its properties at the end of this section.
@@
Let $F$ and $G$ be probability distributions with continuous quantile functions $F^{-1}$ and $G^{-1}$, respectively. Then, for all $p\in[1,\infty)$,
@@ 6 Discussion
As a consequence of our interval-based approach, the proposed decompositions attribute the distance between two distributions $F$ and $G$ entirely to differences in location and dispersion. Of course, distributions are often characterized by additional (higher-order) properties such as skewness or kurtosis attributed to their shape. In our decompositions, differences in higher-order properties sometimes lead to both shift or dispersion components being nonzero simultaneously (e.g., in Example 2.2), but there are no clear-cut connections. The nonparametric nature of our decompositions that aggregate fundamental comparisons of central intervals does not attest to differences in shape. In contrast, decompositions that compare summary statistics as the one in (13), which may do so, provide a rather superficial comparison and are not available for most of the studied distances. Accommodating additional shape components likely requires considerably revised techniques and is reserved for future work.
The decompositions are thus far descriptive tools that appear to be promising both from a theoretical and applied perspective. A next step will be to develop inference techniques that assess the statistical significance of the components in standard one- and two-sample settings as well as more involved settings such as aggregated inflation expectations. While resampling techniques such as permutation tests or the bootstrap are well suited to assess the statistical significance of the overall distance, assessing the significance of individual components is challenging. As the null hypotheses of no difference in shift or location are complex composite hypotheses, simulating them does not seem feasible. A careful look at the sum of the dispersion components of the area validation metric reveals that it can be rewritten as a sum of conditional expectations, which may well facilitate inference for this particular component. We view such developments as a natural avenue for future research.
It remains an open question how the discussed distances can be decomposed for multivariate distributions. For the specific case of the 2-Wasserstein distance, the location component of the moment-based decomposition (13) discussed in Remark 2.4 can be generalized to higher dimensions as follows. If $P$ and $Q$ are probability distributions on $\mathbb{R}^{d}$ with first moments $\mu_{P},\mu_{Q}$, respectively, and finite second moments, the squared (multivariate) 2-Wasserstein distance (Panaretos and Zemel,, 2019, Eq. 1.) can be written as
$\inf_{X\sim P,Y\sim Q}\mathbb{E}\|X-Y\|^{2}=\|\mu_{P}-\mu_{Q}\|^{2}+\inf_{X\sim P,Y\sim Q}\mathbb{E}\|(X-\mu_{P})-(Y-\mu_{Q})\|^{2},$
which may provide a promising starting point for further refinements.
On the other hand, direct generalizations of our nonparametric approach are hindered by the fact that representations of the distances in terms of quantiles are limited to the univariate setting. Our decomposition may however be leveraged in the multivariate setting via a slicing approach; see e.g., Kolouri et al., (2019). Here, a sliced distance is computed as the average distance of linear projections (in all directions) of the distributions to the real line. Accordingly, a sliced decomposition can be obtained by averaging the decomposition components of all linear projections. However, the positive and negative shift components will become equal here, as for each projection, the projection in the opposite direction will lead to an equal contribution to the opposing shift component. Due to the averaging step over all slicing directions, the final shift components will always coincide. One possible improvement would be to fix a hyperplane relative to which the directed components could be defined. Further refinements that incorporate multiple directions in higher dimensions may be feasible. We leave such considerations to future research.
@@ Acknowledgements
We thank Thomas Eife and Thordis Thorarinsdottir for sharing the data sets for the two applications and Christoph Becker, Andreas Eberl, Thomas Eife, Tilmann Gneiting and Melanie Schienle, as well as an anonymous referee, for their insightful comments. T. Dimitriadis and J. Resin (project 502572912) as well as J. Bracher (project 512483310) gratefully acknowledge support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). D. Wolffram was supported by the Helmholtz Association under the joint research school HIDSS4Health – Helmholtz Information and Data Science School for Health. All authors acknowledge the Heidelberg Institute for Theoretical Studies (HITS) as secondary affiliation.
@@
The supplement contains proofs and derivations together with additional illustrations, counterexamples, and details on approximations. Section S1 graphically illustrates the CD decomposition for a range of $\beta$ values. We provide closed-form expressions for the $\operatorname{WD}_{p}$ decomposition in Section S2. Section S3 gives additional counterexamples. Section S4 provides additional details on the computation and approximation of the decompositions. In particular, we provide simple summation formulas for discrete distributions used in Section 5.2 as well as two approximation strategies for settings with limited knowledge of the distributions at hand along with worst case bounds that can be leveraged to derive approximation error rates. Finally, Sections S5S7 contain all proofs of the results presented in Sections 24, respectively.
@@
Consider the distributions $F$ and $G$ that are illustrated in panel (a) of Figure S2. The distribution $F$ places $\frac{1}{4}$ probability mass at 1-1 and $\frac{3}{4}$ probability mass at 0 and $G$ is given by the location-scale transformation $G(x)=F(\frac{x-0.5}{1.8})$. The first three $p$-th power $p$-Wasserstein distances between $F$ and $G$ and their decompositions are
@@
S4 Computation and approximation
When computing the distances and their decompositions numerically, two complications may arise. While in many cases standard numerical integration works fine when computing the required integrals, the nested double integral in the Cramér distance can result in numerical instabilities and long computation times, for example when dealing with discrete distributions. Secondly, the distributions may only be partially specified via a limited number of known points on the distribution or quantile functions, as in the application of Section 5.2.
Here, we provide concrete formulas for discrete distributions in Section S4.1 and outline two approximation strategies for scenarios with incomplete distributions: In Section S4.2, we discuss an approximation based on linear interpolation of the quantile functions, whereas the approximation in Section S4.3 is based on a piecewise constant (nearest neighbor) interpolation. We continue to derive upper and lower worst case bounds for the approximated distances and their decomposition components in Section S4.4. Finally, Section S4.5 illustrates how these bounds can be leveraged for the construction of asymptotic approximation error rates.
As in our applications, we focus on the area validation metric and the Cramér distance in what follows, noting that analogous strategies and results for the higher-order Wasserstein distances are straight-forward generalizations of the case of the AVM.
S4.1 Discrete distributions with finite support
Here, we provide formulas for the AVM and the CD together with their decomposition terms for discrete distributions with finite support. In particular, we consider piecewise constant CDFs $F$ and $G$ that have a finite number of discontinuities (“jumps”) each, which may occur at different quantile levels. Suppose the extended sequence of quantile levels $0=\alpha_{0}<\alpha_{1}<\alpha_{2}<\dots<\alpha_{n}<\alpha_{n+1}=1$ contains all levels at which either quantile function exhibits a jump and, in addition, satisfies $\alpha_{i}=1-\alpha_{n+1-i}$ (symmetry of levels about the median level) with $n$ being odd, i.e., $\alpha_{(n+1)/2}=\frac{1}{2}$. In detail, the set $\{\alpha_{0},\dots,\alpha_{n+1}\}$ is constructed as follows: First, consider the sequence of all quantile levels at which either quantile function exhibits a jump. Second, for each jump level $\alpha$ add the level $1-\alpha$ to the sequence of levels if it is not yet contained in the sequence. Third, add the median level $0.5$ (if not yet contained) as well as the boundary levels 0 and 1 to the sequence. As such, all jump points of $F^{-1}$ and $G^{-1}$ are contained in the ordered sequence $\alpha_{0},\dots,\alpha_{n+1}$, but neither quantile function necessarily jumps at all of these levels.
The distributions $F$ and $G$ are given by piecewise-constant quantile functions [^5] Note that we do not restrict to a particular choice (such as the usual choice of a left-continuous quantile function) of the inherently set-valued quantile at the jump points $\alpha_{i}$, as it does not matter for the integrals in the distances and their decompositions. with
$F^{-1}(\alpha)=f_{i}\quad\text{and}\quad G^{-1}(\alpha)=g_{i}\qquad\text{ if }\alpha\in(\alpha_{i},\alpha_{i+1})$
for $i=0,1,\dots,n$ and some values $f_{0}\leq f_{1}\leq f_{2}\leq\dots\leq f_{n}$ and $g_{0}\leq g_{1}\leq g_{2}\leq\dots\leq g_{n}$.
With $\overline{a}_{i}=\alpha_{i+1}-\alpha_{i}$, the AVM reduces to
$\operatorname{AVM}(F,G)=\sum_{i=0}^{n}\overline{a}_{i}|f_{i}-g_{i}|,$
and the decomposition is given by
$\operatorname{Disp}_{+}^{\operatorname{AVM}}(F,G)$ $=\sum_{i=0}^{\lfloor\frac{n}{2}\rfloor}\overline{a}_{i}\left[(f_{n-i}-g_{n-i})-(f_{i}-g_{i})\right]_{+},$
$\operatorname{Shift}_{+}^{\operatorname{AVM}}(F,G)$ $=2\sum_{i=0}^{\lfloor\frac{n}{2}\rfloor}\overline{a}_{i}\left[\min\{f_{n-i}-g_{n-i},f_{i}-g_{i}\}\right]_{+},$
where the opposing components are obtained via the usual symmetry.
On the other hand, the CD reduces to
$\operatorname{CD}(F,G)=\sum_{i=j}\overline{a}_{i}\overline{a}_{j}|f_{i}-g_{j}|+2\sum_{i<j}\overline{a}_{i}\overline{a}_{j}[f_{i}-g_{j}]_{+}+2\sum_{i>j}\overline{a}_{i}\overline{a}_{j}[g_{j}-f_{i}]_{+}$
with components given by
$\operatorname{Disp}_{+}^{\operatorname{CD}}(F,G)$ $=\sum_{j=0}^{\lfloor\frac{n}{2}\rfloor}\Big{(}\overline{a}_{j}^{2}\left[(f_{n-j}-g_{n-j})-(f_{j}-g_{j})\right]_{+}+2\overline{a}_{j}\sum_{i=j+1}^{\lfloor\frac{n}{2}\rfloor}\overline{a}_{i}\left[(f_{n-i}-g_{n-j})-(f_{i}-g_{j})\right]_{+}\Big{)},$
$\operatorname{Shift}_{+}^{\operatorname{CD}}(F,G)$ $=\sum_{i=0}^{\lfloor\frac{n}{2}\rfloor}\sum_{j=0}^{\lfloor\frac{n}{2}\rfloor}2\overline{a}_{i}\overline{a}_{j}\left(\left[\min\{f_{n-i}-g_{n-j},f_{i}-g_{j}\}\right]_{+}+\left[f_{i}-g_{n-j}\right]_{+}\right).$
S4.2 Piecewise linear approximation
Here, we outline an approximation based on a linear interpolation of the quantile functions that can be applied when the distributions $F$ and $G$ are only partially known via a limited number of points on the distribution or quantile functions. For instance in the application from Section 5.2, the assigned bin probabilities together with the bin boundaries (i.e., the values separating the bins) identify the distribution and quantile functions at these boundary values. However, no further information on the quantile functions can be inferred from the responses.
In Figure 13 of the main text, we illustrate the points where the distributions are known by the solid points in the quantile spread plots. To compute the Cramér and Wasserstein distances and their associated decompositions in this case, the quantile functions need to be approximated. For this, we adopt a simple linear interpolation between the known quantiles of the distributions as illustrated in Figure 13 with the straight lines connecting the points. In terms of the underlying histogram forecasts, this corresponds to uniformly distributing the probability mass within each bin.
S4.3 Piecewise constant approximation
Here, we outline a different approximation method that relies on a piecewise constant (nearest neighbor) interpolation of the quantile functions. This approximation may be of use if both distributions are only known via the quantile values at a fixed (and typically identical) set of quantile levels. This is a common way of storing predictive distributions in a parsimonious format; see e.g., Cramer et al., (2022).
As an alternative to the linear interpolation from the previous section, we propose to approximate the unknown quantile functions at each level via the closest known quantile value of the distributions here. The resulting approximate quantile functions give rise to discrete distributions, to which the formulas from Section S4.1 can be applied. In some cases, applying this approach on a fine grid of levels may also help to avoid numerical instabilities or improve computational efficiency, especially for the CD.
Suppose the values of the quantile functions $F^{-1}$ and $G^{-1}$ are known at levels $0\leq\beta_{1}<\dots<\beta_{K}\leq 1$ and $0\leq\gamma_{1}<\ldots<\gamma_{L}\leq 1$, respectively. Let $\beta_{0}=-\beta_{1},\beta_{K+1}=2-\beta_{K},\gamma_{0}=-\gamma_{1},\gamma_{L+1}=2-\gamma_{L}$. We define two approximate discrete distributions $\widetilde{F},\widetilde{G}$ with quantile functions given by
$\widetilde{F}^{-1}(\beta)$ $=F^{-1}(\beta_{k})\quad\text{for}\quad\beta\in\left(\frac{\beta_{k}+\beta_{k-1}}{2},\frac{\beta_{k}+\beta_{k+1}}{2}\right),\qquad\text{and}$
$\widetilde{G}^{-1}(\gamma)$ $=G^{-1}(\gamma_{\ell})\quad\text{for}\quad\gamma\in\left(\frac{\gamma_{\ell}+\gamma_{\ell-1}}{2},\frac{\gamma_{\ell}+\gamma_{\ell+1}}{2}\right).$
Then, the distances and their decompositions can be approximated by applying the formulas provided in Section S4.1 to the approximate distributions $\widetilde{F}$ and $\widetilde{G}$. In particular, we propose the approximations